# Radial Fast Entangling Gates Under Micromotion in Trapped-Ion Quantum Computers

### v1

@@ II Background Theory: Quantum dynamics in Radio-frequency traps
where $m$ is the mass of the ion, $\Omega_{\rm RF}$ is the frequency of the RF drive with phase $\phi_{\rm RF}$ at $t=0$. The parameters $a_{x}$ and $q_{x}$ are trapping parameters proportional to the strength of the DC and AC components of the electric potential, respectively [Leibfried2003b].
In our model of the dynamics, we assume the ions are fixed in the $y$ axis to $y^{(1)}_{0}=y^{(2)}_{0}=0$ and are in their axial equilibrium positions: $z^{(1)}_{0}=-z^{(2)}_{0}=d/2$, with $d$ the on-axis spacing between ions as shown in Fig. 1(a). While mathematically convenient, this assumption is not required for the fast gate mechanism, which as a geometric phase gate protocol is insensitive to the precise initial conditions of the ions’ motion, e.g. due to thermal fluctuations [Garcia-Ripoll2003c]. The potential energy describing the 1D dynamics of the system along the radial ($x$) axis is then given by:
$V(x^{(1)},x^{(2)},t)=\sum_{j=1,2}V_{\rm RF}(x^{(j)},t)+\frac{e^{2}}{4\pi\epsilon_{0}}\frac{1}{\sqrt{(x^{(1)}-x^{(2)})^{2}+d^{2}}},$ (2)
@@
where $e$ is the elementary charge. This potential has a stationary point at $x^{(1)}=x^{(2)}=0$, which corresponds to ion equilibria along the null of the RF trap. Assuming small excursions in phase space, such that we can expand the Coulomb term to quadratic order in $(x^{(1)}-x^{(2)})/d$, and transforming into centre-of-mass (CM) and relative/breathing (BR) coordinates, i.e. $x_{\rm CM/BR}=(x^{(1)}\pm x^{(2)})/\sqrt{2}$, the potential energy of the system may be expressed as:
$V_{\rm RF}(x_{\rm CM},x_{\rm BR})$ $\approx\frac{1}{2}m\sum_{\alpha={\rm CM,BR}}\lambda_{\alpha}(t)x_{\alpha}^{2}$ (3)
@@
where $\lambda_{\alpha}(t)\equiv(\Omega_{\rm RF}/2)^{2}[a_{\alpha}-2q_{x}\cos(\Omega_{\rm RF}t+\phi_{\rm RF})]$ in terms of the mode-dependent trapping parameter $a_{\alpha}=a_{x}-4\Gamma_{\alpha}(d)/(m\Omega_{\rm{RF}}^{2})$. In Appendix B we show that for a one-dimensional system of $N=2$ ions in a common harmonic potential, $\Gamma_{\rm CM}(d)=0$ and
$\Gamma_{\rm BR}=\frac{e^{2}}{4\pi\epsilon_{0}d^{3}}\,.$ (4)
@@
Here we adopt a Heisenberg picture approach [Leibfried2003b, Ji1995, Wu2021], where the position and momentum operators of each motional mode evolve according to the equations of motion:
@@
These equations of motion admit the following solution [Leibfried2003b]:
$\hat{x}_{\alpha}(t)$ $=\sqrt{\frac{\hbar}{2m\omega_{\alpha}}}\left(\hat{a}_{p}u^{*}_{\alpha}(t)+\hat{a}^{\dagger}_{\alpha}u_{\alpha}(t)\right)\,$, (7a)
@@
which is a solution to the Mathieu-Hill equation, $\ddot{u}_{\alpha}(t)=-\lambda_{\alpha}(t)u_{\alpha}(t)$. These mode functions include a low frequency component, which we refer to as the secular mode frequency $\omega_{\alpha}\equiv\beta_{\alpha}\Omega_{\rm RF}/2$  [^1] This definition of the secular frequency is subtly different from the definition in terms of the Wronskian identity $u^{*}_{\alpha}(t)\dot{u}_{\alpha}(t)-u_{\alpha}(t)\dot{u}_{\alpha}^{*}(t)=2i\omega_{\alpha}$ [Leibfried2003b]., and high-frequency components at integer multiples of the RF drive – i.e. micromotion. The Fourier coefficients in Equation (8) satisfy the recurrence relation
@@
$C_{\alpha,n}$ $=\frac{C_{\alpha,n+1}}{D_{\alpha,n}-\frac{1}{D_{\alpha,n-1}-\frac{1}{D_{\alpha,n-2}-\dots}}}=\frac{C_{\alpha,n-1}}{D_{\alpha,n}-\frac{1}{D_{\alpha,n+1}-\frac{1}{D_{\alpha,n+2}-\dots}}}\,.$ (10)

### v2

@@ Appendix A Quantum dynamics in radio-frequency traps
where $m$ is the mass of the ion, $\Omega_{\rm RF}$ is the frequency of the RF drive with phase $\phi_{\rm RF}$ at $t=0$. The parameters $a_{x}$ and $q_{x}$ are trapping parameters proportional to the strength of the DC and AC trapping fields, respectively [35].
Then, assuming two ions fixed to their equilibrium positions in the $y$ and $z$ axes, – $y^{(1)}_{0}=y^{(2)}_{0}=0$ and $z^{(1)}_{0}=-z^{(2)}_{0}=d/2$, the potential energy of the two-ion system along the radial ($x$) axis is given by:
$V(x^{(1)},x^{(2)},t)=$ $\sum_{j=1,2}V_{\rm RF}(x^{(j)},t)$ (16)
$+\frac{e^{2}}{4\pi\epsilon_{0}}\frac{1}{\sqrt{(x^{(1)}-x^{(2)})^{2}+d^{2}}},$
@@
where $e$ is the elementary charge. This potential has a stationary point at $x^{(1)}=x^{(2)}=0$, which corresponds to ion equilibria along the null of the RF trap (see Fig. 1). Assuming small excursions in phase space, such that we can expand the Coulomb term to quadratic order in $(x^{(1)}-x^{(2)})/d$, and transforming into centre-of-mass (CM) and relative/stretch (SR) coordinates, i.e. $x_{\rm CM/SR}=(x^{(1)}\pm x^{(2)})/\sqrt{2}$, the potential energy of the system may be expressed as:
$V_{\rm RF}(x_{\rm CM},x_{\rm SR})$ $\approx\frac{1}{2}m\sum_{\alpha={\rm CM,SR}}\lambda_{\alpha}(t)x_{\alpha}^{2}$ (17)
@@
where $\lambda_{\alpha}(t)\equiv(\Omega_{\rm RF}/2)^{2}[a_{\alpha}-2q_{x}\cos(\Omega_{\rm RF}t+\phi_{\rm RF})]$ in terms of the mode-dependent trapping parameter $a_{\alpha}=a_{x}-4\Gamma_{\alpha}(d)/(m\Omega_{\rm{RF}}^{2})$. For a one-dimensional system of $N=2$ ions in a common harmonic potential, $\Gamma_{\rm CM}(d)=0$ and
$\Gamma_{\rm SR}=\frac{e^{2}}{4\pi\epsilon_{0}d^{3}}\,.$ (18)
@@
Here we adopt a Heisenberg picture approach [35, 29, 77], where the position and momentum operators of each motional mode evolve according to the equations of motion:
@@
These equations of motion admit the following solution [35]:
$\hat{x}_{\alpha}(t)$ $=\sqrt{\frac{\hbar}{2m\omega_{\alpha}}}\left(\hat{a}_{\alpha}u^{*}_{\alpha}(t)+\hat{a}^{\dagger}_{\alpha}u_{\alpha}(t)\right)\,$, (21a)
@@
which is a solution to the Mathieu-Hill equation, $\ddot{u}_{\alpha}(t)=-\lambda_{\alpha}(t)u_{\alpha}(t)$. These mode functions include a low frequency component, which we refer to as the secular mode frequency $\omega_{\alpha}\equiv\beta_{\alpha}\Omega_{\rm RF}/2$, and high-frequency components at integer multiples of the RF drive – i.e. micromotion. The Fourier coefficients in Equation (22) satisfy the recurrence relation
@@
$C_{\alpha,n}$ $=\frac{C_{\alpha,n+1}}{D_{\alpha,n}-\frac{1}{D_{\alpha,n-1}-\frac{1}{D_{\alpha,n-2}-\dots}}}$ (24)
$=\frac{C_{\alpha,n-1}}{D_{\alpha,n}-\frac{1}{D_{\alpha,n+1}-\frac{1}{D_{\alpha,n+2}-\dots}}}\,.$